This procedure is not recommended! You'll see why after you work the problem. You are on an aluminum ladder that is standing on the ground, trying to fix an electrical connection with a metal screwdriver having a metal handle. Your body is wet because you are sweating from the exertion; therefore, it has a resistance of 1.80 kΩ .

Part A If you accidentally touch the "hot" wire connected to the 120 V line, how much current will pass through your body? I =

Part B Is this amount enough to be dangerous? (The maximum safe current is about 5.00 mA .)

Part C How much electrical power is delivered to your body?
Answer #1

from ohms law

I = V/R

V = potential


R = resistance


I = 120/(1.8*10^3) = 0.067 A = 67 mA


---------------

part(B)

YES dangerous

--------------

part B


power = I^2*R


P = (67*10^-3)^2*1.8*10^3 = 8.1 W